9. The current barrel of oil price is equal to 47.23 USD/barrel. We expect that the price will decrease to 46.89 USD/barrel. The bank offers call and put option with strike price equal to 47.07 USD/barrel. The option premium is equal to 0.12 USD/barrel. The one option contract contains 100 barrels. Investor has 6 000 USD. How many and what options will you purchase? What will be result of the operation if your expectations will prove to be true?
Answer #1
1) As the expected future price per barrel of $46.89 is less
than the strike price-option premium (per barrel) of
$47.07-$0.12 = $46.95, it would be profitable to buy
put options.
On the expiry date the oil can be purchased from the
open market at $46.89 per barrel and then can be sold
against the put optio at $47.07 per barrel.
The gain per barrel = 47.07-0.12-46.89 = $0.06 per barrel.
2) Number of put option contracts that can be purchased = 6000/(0.12*100) = 500 put options
(500 contracts is equal to = 500*100 = 50000 barrels)
3) Purchase cost of 50000 barrels from the open market = 46.89*50000 = $ 23,44,500
Option premium paid = 50000*0.12 = $           6,000
Receipts for sale under put option = 50000*47.07 = $ 23,53,500
Profit from the operation = 2353500-2344500-6000 = $           3,000
[Alternatively 50000*0.06 = $3000]